# Polarized multiwavelength emission from pulsar wind – accretion disk interaction in a transitional millisecond pulsar

### v1

@@ Abstract
Transitional millisecond pulsars (tMSPs) bridge the evolutionary gap between accreting neutron stars in low-mass X-ray binaries and millisecond radio pulsars. These systems exhibit a unique subluminous X-ray state characterized by the presence of an accretion disk and rapid switches between high and low X-ray emission modes. The high mode features coherent millisecond pulsations spanning from the X-ray to the optical band. We present multiwavelength polarimetric observations of the tMSP PSR J10230038 aimed at conclusively identifying the physical mechanism powering its emission in the subluminous X-ray state. During the high mode, we detect polarized emission in the 2–6 keV energy range, with a polarization degree of and a polarization angle of (1) measured counterclockwise from the North celestial pole towards East. At optical wavelengths, we find a polarization degree of and a polarization angle aligned with that in the soft X-rays, suggesting a common physical mechanism operating across these bands. Remarkably, the polarized flux spectrum matches the pulsed emission spectrum from optical to X-rays. The polarization properties differ markedly from those observed in other accreting neutron stars and isolated rotation-powered pulsars and are also inconsistent with an origin in a compact jet. Our results provide direct evidence that the polarized and pulsed emissions both originate from synchrotron radiation at the shock formed where the pulsar wind interacts with the inner regions of the accretion disk.
@@ 1 Introduction
The evolution of accreting neutron stars (NSs) in low-mass X-ray binaries (LMXBs) as progenitors of millisecond radio pulsars (MSPs) involves distinct phases [1], with transitional MSPs (tMSPs) representing one of the most intriguing stages of this process due to their unique properties [23, 95]. PSR J10230038 (hereafter J1023) is the prototypical tMSP [4], offering an exceptional opportunity to study these enigmatic objects in detail. Detected in 2007 as a radio MSP with a spin period of 1.69 ms, J1023 orbits a 0.24 star every 4.75 hours [4]. In 2013, J1023 experienced dramatic changes in its emission properties, ceasing its radio pulsations and exhibiting increased emission across the electromagnetic spectrum [112, 101]. J1023 has since remained in an active state, characterized by subluminous X-ray emission with an average luminosity of erg s-1 in the 0.3–80 keV energy range [33] – over four orders of magnitude below the Eddington luminosity for a NS. This emission alternates between high-intensity (70–80% of the time) and low-intensity (20–30%) modes, with occasional brief flares reaching erg s-1 [96, 81, 5, 18]. Mode switches occur within 10–30 s, with low modes lasting from tens of seconds to minutes. During the high mode, J1023 exhibits coherent pulsed emission at the pulsar spin period at X-ray, ultraviolet (UV), and optical wavelengths, pointing to processes linked to the NS rotation [5, 98, 90, 70]. This pulsed emission disappears during the low mode. The UV, optical, and near-infrared emissions mainly originate from an accretion disk surrounding the NS and the irradiated companion star, leading to flickering and flaring [74, 97, 59, 109]. Bright radio and mm continuum emission is observed, with increased radio emission during the low mode and brief mm-band flares during switches from high to low modes [37, 19, 10].
@@
The unique behavior of J1023 raises fundamental questions about the mechanisms powering its emission in the active state. J1023 releases rotational kinetic energy at a rate of erg s-1 [3], about six times its average X-ray luminosity. This has sparked debates on whether the emission of tMSPs like J1023 in this active state is powered by pulsar rotation, mass accretion, or a combination of both [19, 98, 122, 82]. Multiband polarization measurements provide a key diagnostic tool for investigating these processes and solving this enigma.
@@
This paper presents the results from the first multiwavelength polarimetric campaign of J1023 and is structured as follows. Section 2 describes the observations and data processing techniques employed across the X-ray, optical, and radio bands using the Imaging X-ray Polarimetry Explorer (IXPE), the Neutron Star Interior Composition Explorer Mission (NICER), the Neil Gehrels Swift Observatory (Swift), the Very Large Telescope (VLT), and the Karl G. Jansky Very Large Array (VLA). Section 3 details the data analysis and results, encompassing the selection of emission modes and timing analysis of the IXPE data, and the polarimetric analysis across different wavelengths. We also present the spectral energy distributions (SEDs) derived from our observations. Section 4 discusses the implications of our findings on the emission mechanisms powering J1023 during its active state, evaluating scenarios such as accretion-powered processes, rotation-powered mechanisms analogous to isolated pulsars, and interactions between the pulsar wind and the accretion flow. Section 5 summarizes how our multiwavelength polarimetric observations provide critical insights into the enigmatic behavior of tMSPs like J1023.

### v2

@@ Abstract
Transitional millisecond pulsars (tMSPs) bridge the evolutionary gap between accreting neutron stars in low-mass X-ray binaries and millisecond radio pulsars. These systems exhibit a unique subluminous X-ray state characterized by the presence of an accretion disk and rapid switches between high and low X-ray emission modes. The high mode features coherent millisecond pulsations spanning from the X-ray to the optical band. We present multiwavelength polarimetric observations of the tMSP PSR J10230038 aimed at conclusively identifying the physical mechanism powering its emission in the subluminous X-ray state. During the high mode, we report a probable detection of polarized emission in the 2–6 keV energy range, with a polarization degree of and a polarization angle of measured counterclockwise from the North celestial pole towards the East (99.7% c.l.; uncertainties are quoted at 1). At optical wavelengths, we find a polarization degree of and a polarization angle aligned with that in the X-rays, suggesting a common physical mechanism operating across these bands. Remarkably, the polarized flux spectrum matches the pulsed emission spectrum from optical to X-rays. The polarization properties differ markedly from those observed in other accreting neutron stars and isolated rotation-powered pulsars and are also inconsistent with an origin in a compact jet. Our results provide direct evidence that the polarized and pulsed emissions both originate from synchrotron radiation at the boundary region formed where the pulsar wind interacts with the inner regions of the accretion disk.
show]maria.baglio@inaf.it ††thanks: The first two authors equally contributed to this work.
show]cotizelati@ice.csic.es ††thanks: The first two authors equally contributed to this work.
@@ 1 Introduction
The evolution of accreting neutron stars (NSs) in low-mass X-ray binaries (LMXBs) as progenitors of millisecond radio pulsars (MSPs) involves distinct phases (M. A. Alpar et al., 1982), with transitional MSPs (tMSPs) representing the most intriguing stage of this process due to their unique properties (S. Campana & T. Di Salvo, 2018; A. Papitto & D. de Martino, 2022). PSR J10230038 (J1023) is the prototypical tMSP (A. M. Archibald et al., 2009), offering an opportunity to study these enigmatic objects in detail. Detected in 2007 as a radio MSP with a spin period of 1.69 ms, J1023 orbits a 0.24 star every 4.75 hours (A. M. Archibald et al., 2009).
@@
In 2013, J1023 experienced dramatic changes in its emission properties, ceasing its radio pulsations and exhibiting increased emission across the electromagnetic spectrum (B. W. Stappers et al., 2014; A. Patruno et al., 2014). J1023 has since remained in an active state, characterized by subluminous X-ray emission with an average luminosity of erg s-1 in the 0.3–80 keV energy range (F. Coti Zelati et al., 2018) – over four orders of magnitude below the Eddington luminosity for a NS. This emission alternates between high-intensity (70–80% of the time) and low-intensity (20–30%) modes, with occasional brief flares reaching 1035 erg s-1 (A. Papitto et al., 2013; M. Linares, 2014; A. M. Archibald et al., 2015; S. Bogdanov et al., 2015). Mode switches occur within 10–30 s, with low modes lasting from tens of seconds to minutes. During high modes, which typically last from minutes to several tens of minutes, J1023 exhibits pulsed emission at the pulsar spin period at X-ray, ultraviolet (UV), and optical wavelengths, pointing to processes linked to the NS rotation (A. M. Archibald et al., 2015; A. Papitto et al., 2019; A. Miraval Zanon et al., 2022; G. Illiano et al., 2023).
This pulsed emission disappears during the low mode. The UV, optical, and near-infrared emissions mainly originate from an accretion disk surrounding the NS and the irradiated companion star, leading to flickering and flaring (M. R. Kennedy et al., 2018; A. Papitto et al., 2018; P. Hakala & J. J. E. Kajava, 2018; T. Shahbaz et al., 2018). Bright radio–mm continuum emission is observed, with increased radio emission during the low mode and brief mm-band flares during switches from high to low modes (A. T. Deller et al., 2015; S. Bogdanov et al., 2018; M. C. Baglio et al., 2023).
@@
The behavior of J1023 raises fundamental questions about the mechanisms powering its emission in the active state. J1023 releases rotational kinetic energy at a rate of erg s-1 (A. M. Archibald et al., 2013), about six times its average X-ray luminosity. This has sparked debates on whether the emission of tMSPs like J1023 in this active state is powered by pulsar rotation, mass accretion, or a combination of both (S. Bogdanov et al., 2018; A. Papitto et al., 2019; A. Veledina et al., 2019; M. Linares et al., 2022). Multiband polarization measurements provide a diagnostic tool for investigating these processes and solving this enigma.
@@
This Letter presents the results from the first multiwavelength polarimetric campaign of J1023 and is structured as follows. Section 2 describes the observations and data processing techniques employed across the X-ray, optical, and radio bands using the Imaging X-ray Polarimetry Explorer (IXPE), the Neutron Star Interior Composition Explorer Mission (NICER), the Neil Gehrels Swift Observatory (Swift), the Very Large Telescope (VLT), and the Karl G. Jansky Very Large Array (VLA). Section 3 details the data analysis and results, encompassing the selection of emission modes and timing analysis of the IXPE data, and the polarimetric analysis across different wavelengths. We also present the spectral energy distributions (SEDs) derived from our observations. Section 4 discusses the implications of our findings on the emission mechanisms powering J1023 during its active state, evaluating scenarios such as accretion-powered processes, rotation-powered mechanisms analogous to isolated pulsars, and interactions between the pulsar wind and the accretion flow. Section 5 summarizes how our multiwavelength polarimetric observations provide critical insights into the enigmatic behavior of tMSPs like J1023.